High-Low Methed Diamond Inc, has decided to use the high-low method to estimate the total cost and the fixed and variable cost components af the total cost. The dats for various levels o production are as follows: Units Produced 12,000 15,000 21,000 Total Costs $424,000 496,000 640,000 a. Determine the variable cost per unit and the total fixed cost Variable cost (Round to two decimal places.) Total fixed cost b. Based on part (a), estimate the total cost for 17,000 units of production. Total cost for 17,000 units 4 per unit Omds the aifferance e betwsan the highest and ioweet total costs by te diference batween the highost and lowest production unts ths varable cost trom the morths total cost to amve at tha d cost. The resuting amount s Mutply the vanabie unt cost by The numbar ofunts for should se the sane for ethar the highast or lowast production month Next All work saved
Answer #1

Cost

No. of activities

A

High Level

$     6,40,000.00

21000

B

Low Level

$    4,24,000.00

12000

C=A-B

Difference

$     2,16,000.00

9000

A

Cost difference

$     2,16,000.00

B

No. of Activities difference

9000

C=A/B

Variable cost per unit

$                 24.00

High Level

Low Level

A

No. of activities

21000

12000

B

Variable cost per unit

$                 24.00

$             24.00

C=AxB

Total Variable cost

$     5,04,000.00

$ 2,88,000.00

D

Total cost

$     6,40,000.00

$ 4,24,000.00

E=D-C

Total fixed cost

$     1,36,000.00

$ 1,36,000.00

Requirement (a)

Total Fixed cost = $136,000

.

Requirement (b)

Cost at 17000 units

A

Units

17000

B

Variable cost per unit

$                 24.00

C=AxB

Total Variable cost

$     4,08,000.00

D

Fixed cost

$     1,36,000.00

E=D+C

Total cost

$     5,44,000.00

Total cost at 17000 units = $544000